The following charges are located inside a metallic container on unknown shape: -4.10 ?C, -5.10 ?C, 65.0 ?C, and 58.0 ?C. Calculate the net electric flux, in Nm2/C, through the container.
Answer #1

he total of the electric flux out of a closed surface is equal to the charge enclosed divided by thepermittivity.

The sum of the flux i proportional to the total charge enclosed. electric c ДА perpendicular 0

Total Charge Enclosed = (- 4.10 - 5.10 + 65.0 + 58.0)*10-6 C = 113.8*10-6 C

Total Flux = (113.8*10-6)/(8.85*10-12) = 12.859*106 Nm2/C